# Optimizing Language Models for Human Preferences is a Causal Inference Problem

### v1

@@ 1 Introduction
The ability to learn human preferences from direct outcome data significantly broadens the scope of problems that can be addressed by learning from human feedback. Consider the task of inducing a language model to unlearn hate speech. Current unlearning approaches typically use paired data in the format ([hate speech], [alternative text]), with the preferred text being the latter. Constructing alternative texts can be difficult [Eldan and Russinovich, 2023, Maini et al., 2024], and rather than fully unlearning the hate speech, the language model is instead trained to preferentially generate the alternative text [Patil et al., 2024]. Direct outcome data, by contrast, allows texts to be directly marked as hateful and removed from the language model without learning or requiring an alternative text.
@@
Therefore, we posit that language model optimization should be viewed as a causal problem in order to ensure that the optimal language model causes preferred outcomes. In this paper, we introduce a causal formulation of the language model optimization problem. The solution to this optimization problem finds how to intervene on the text distribution of the generating model to best cause an optimal outcome (i.e., generation of human-preferred texts).
We observe that in the direct outcome setting, it is possible in practice to guarantee that the observed relationship between the text and the outcome is causal by leveraging crowdsourced datasets. Due to random assignment of texts to readers, crowdsourced datasets are not subject to external confounding and can in fact be viewed as randomized experiments [Lin et al., 2023]. Building on this observation, we present two methodological contributions that enable causal language model optimization on direct outcome datasets. First, we develop causal preference optimization (CPO), which solves an unbiased surrogate objective for the causal optimization problem. Next, we extend this to doubly robust CPO (DR-CPO), which improves on CPO by reducing the variance of the surrogate objective via outcome modeling while retaining provably strong guarantees on bias.
@@ 3 A Causal View of Language Model Optimization
When a language model $f$ is trained or fine-tuned to generate texts that are consistent with human preferences, the implicit goal can be seen as optimizing texts $X\sim P^{f}$—texts generated from the model ff—with respect to some outcome $Y$. In this paper, we consider a direct outcome data format $\mathcal{D}_{O}=\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\}$, where $X_{i}$ is a text that individual $i$ interacts with and $Y_{i}$ is any numerical response of the individual to the texts (e.g., ratings, either binary or scalar).
@@
These hypothetical outcomes can be formalized using the potential outcomes framework [Neyman, 1923 [1990], Rubin, 1974]: over text space $\mathcal{X}$, for each individual $i$ we posit the existence of a potential outcome function $Y_{i}:\mathcal{X}\to\mathbbm{R}$, where $Y_{i}(x)$ encodes their potential real-valued response if given text xx. [^1] This notation implicitly rules out the possibility that an individual’s responses can be affected by the texts given to others—a common assumption in causal inference [Rubin, 1974]. We emphasize that most individuals’ potential outcomes are not observed, and so $Y_{i}(x)$ denotes the response individual $i$ would have had had they seen text $x$, possibly contrary to reality. This is also commonly known as the counterfactual.
@@
$\underset{f}{\arg\max}\;V(f)=\underset{f}{\arg\max}\;\mathbb{E}_{X\sim P^{f}}[\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}[Y(X)]]$ (3)
@@
Reframing our optimization problem as a causal inference problem allows us to draw on solutions from statistical causal inference—in particular, the use of randomized experiments to identify causal effects and approximate observing all potential outcomes. [^2] As discussed above, crowdsourced datasets are in fact randomized experiments, since annotators are randomly assigned to read texts. We formalize such a randomized experiment and/or crowdsourced annotated dataset as $\mathcal{D}_{R}=\{(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\}$ where texts $X_{i}$ are drawn i.i.d. from a randomization distribution $P^{R}$ and individuals with potential outcome functions $Y_{i}(\cdot)$ are drawn i.i.d. from the population $\mathcal{G}$. This induces a distribution on the observed responses $Y_{i}=Y_{i}(X_{i})$ that we denote as $P^{R}_{y}$.
@@ 4.1 Causal Preference Optimization
[…]
@@ 4.2 Doubly Robust Causal Preference Optimization
An importance weighting estimator like the CPO value function is a natural solution for estimating causal quantities when randomized experimental data such as crowdsourced data is available. However, CPO optimizes over only the experimental data, and so it can be further improved by the addition of an outcome modeling term that predicts outcomes on unlabeled texts. The combination of IPW and outcome modeling yields a doubly robust estimator (DR-CPO) that reduces the variance of CPO and improves its generality while still remaining unbiased for the true causal optimization problem.
@@
Importantly, because $P^{R}$ is known in randomized experiments, including the crowdsourced data setting, it does not need to be estimated, which means that condition (1) is always fulfilled for DR-CPO. Therefore, $\widehat{V}_{DR}(f)$ is guaranteed to be unbiased for $V_{DR}$ even if the outcome model is incorrect. In other words, DR-CPO is robust to misspecification of $\widehat{g}$, as the IPW term in its value function corrects for any bias from the predicted outcomes.
[…]
@@
[…]
@@
Condition (i) limits the component of the variance difference that arises due to Monte Carlo error from taking $m$ samples from the reference language model $f^{0}$. Then the main comparison is the difference between the variance of the expected outcome $g(x)$, and the variance of the prediction error for the model. As a result, under condition (ii), we can expect the variance of $\widehat{V}_{DR}$ to be lower than the variance of $\widehat{V}_{IPW}$ (e.g., if $(\widehat{g}(x)-g(x))^{2}<g(x)^{2}$).
We note that in a different data setting where the true $P^{R}$ is unknown, a well-estimated $\widehat{g}$ that is close to the true $g$ can also help bias-correct any mis-estimation of $\widehat{P}^{R}(X)$. This may occur, for instance, when no text experiment or crowdsourced dataset is available, but a large amount of clean data exists to train the outcome model.
@@ 4.3 Relationship to Existing Approaches
First, notice that the outcome modeling term $V_{out}$ is itself a way of identifying $V(f)$ from the observed data. Outcome modeling relies entirely on the predictive model $\widehat{g}$ and unlabeled texts $\widetilde{X}\sim P^{f^{0}}$ and therefore does not require an experimental dataset $\mathcal{D}_{R}$. However, if $\widehat{g}(X)$ is not a good outcome model, then $\widehat{V}_{out}(f)$ will be biased with respect to $V(f)$. In particular, reward models trained on confounded data may be misspecified, since they can only capture the relationship between the text and the response—and not the confounders that have additionally influenced the response. While these issues are remedied if the confounding is also fully modeled, confounders are extremely difficult to measure fully in text data.
Optimization of $V_{out}(f)$ is closely related to RLHF via proximal policy optimization (PPO) in the direct outcome data setting. In particular, the PPO policy loss term can be seen as a version of outcome modeling in which the reward model is trained on paired completion data. Under these conditions, the RLHF reward model is analogous to the outcome model $g$, and the RLHF policy loss is mathematically equivalent to $V_{out}(f)$ (details in Appendix B).
Likewise, DPO shares similarities with CPO. Both DPO and CPO fine-tune a language model for human feedback by directly using a preference dataset rather than relying on reward or outcome modeling. The DPO objective is similar to $V_{IPW}$ in that it directly increases the likelihood of texts corresponding to desired outcomes through importance weighting. However, because of the paired nature of the data, DPO increases the density ratio between preferred and non-preferred examples, while CPO directly increases the probability of texts with desired outcomes and decreases the probability of texts with non-desired outcomes. Given a paired data setting, the DPO objective could possibly be recovered from $V_{IPW}$; we leave this derivation for future work.
@@ 5.1 Datasets
Hate Speech (binary outcome). The Hate Speech dataset [Qian et al., 2019] consists of comments from the social media sites Reddit and Gab. Outcomes are collected via crowdsourcing and indicate whether the annotator percieves the comment to be hate speech. The Reddit comments are chosen from subreddits where hate speech is more common, and Gab is a platform where users sometimes migrate after being blocked from other social media sites. The optimization goal for this dataset is to generate texts that are less hateful on average.
@@
Confounded (scalar outcome). The Confounded dataset is a version of the Hong Kong dataset where we have induced confounding. We consider the strongest possible form of confounding: the confounder is fully correlated with the outcome, resulting in all outcomes being negations of the original outcomes. This dataset is used to train outcome models. We include this dataset with the realistic expectation that text data is often confounded, which poses a threat to outcome model-based approaches. Therefore, it is necessary to evaluate how different optimization approaches fare under confounding. Like the Hong Kong dataset, the optimization goal for this dataset is to generate texts with high outcomes on average.
@@ 5.2 Implementation
Evaluation. To evaluate how well optimization for human preferences has occurred, we use a text preference framework in which a reader is asked to choose the better (with respect to the outcome) of a pair of texts generated by two different methods. Using GPT-4 as a proxy for human annotators, we compare pairs of (method, baseline) completions for the same prompt; across all pairs, we compute method win rates and compute 95% confidence intervals. Since the datasets used for these experiments contain one text per sample rather than a prompt and a completion, we create prompts on the evaluation set by truncating each text to a random length.
@@
Methods. We evaluate language models optimized using CPO and DR-CPO. As our baselines, we consider language models that have been fine-tuned on texts from each of the task datasets (FT), as well as models optimized using the outcome modeling value function $V_{out}(f)$. Since—as we discuss in Section 4.3—the $V_{out}(f)$ objective is mathematically equivalent to the RLHF objective, we refer to this baseline as $O$-RLHF (offline outcome RLHF).
We use Llama 2 7B [Touvron et al., 2023] as our base language model and fine-tune with low-rank adaptation [Hu et al., 2022]. All optimizations are applied after fine-tuning on text from the task dataset.
Choice of $f^{0}$. When optimizing with DR-CPO or $O$-RLHF, any generative language model may be used as $f^{0}$, the fixed language model from which texts are drawn as input to the outcome model. One key consideration is whether $f^{0}$ should be a pre-trained model or whether it should be a model that has been fine-tuned on text relevant to the task—for instance, the randomized experiment dataset $\mathcal{D}_{R}$.
In practice, we choose a pre-trained model as $f^{0}$ to leverage the diversity of texts such models tend to generate. If $\widehat{g}(X)$ is a good outcome model, then predicted outcomes on these texts will still be close to the true outcomes, and DR-CPO and $O$-RLHF will benefit from outcome modeling.
Choice of $\widehat{P}^{R}$. In Section 4.1, we mention briefly that although the distribution of texts $P^{R}(X)$ under the randomized experiment is known, it can be helpful empirically to instead compute an estimated $\widehat{P}^{R}(X)$. This is generally due to the fact that the sample probability of each text $X$ may not actually be equal to its theoretical probability merely by chance [Hájek, 1971, Särndal et al., 2003].

### v2

@@ 1 Introduction
The ability to learn human preferences from direct outcome data significantly broadens the scope of problems that can be addressed by learning from human feedback. Consider the task of inducing a language model to unlearn hate speech. Current unlearning approaches typically use paired data in the format ([hate speech], [alternative text]), with the preferred text being the latter. Constructing alternative texts can be difficult [Eldan and Russinovich, 2023, Maini et al., 2024], and rather than fully unlearning the hate speech, the language model is instead trained to preferentially generate the alternative text [Patil et al., 2024]. Direct outcome data, by contrast, allows texts to be directly marked as hateful and removed from the language model without learning or requiring an alternative text. Moreover, direct outcome datasets greatly outnumber paired completion datasets in NLP: at the time of writing, for instance, the HuggingFace Datasets hub contains thousands of direct outcome datasets and fewer than 50 paired completion datasets.
@@
Therefore, we posit that language model optimization should be viewed as a causal problem in order to ensure that the optimal language model causes preferred outcomes. In this paper, we introduce a novel causal formulation of the language model optimization problem. The solution to this optimization problem finds how to intervene on the text distribution of the generating model to best cause an optimal outcome (i.e., generation of human-preferred texts).
We observe that in the direct outcome setting, it is possible in practice to guarantee that the observed relationship between the text and the outcome is causal by leveraging crowdsourced datasets. Due to random assignment of texts to readers, crowdsourced datasets are not subject to external confounding and can in fact be viewed as randomized experiments [Lin et al., 2023]. Building on this observation, we present two methodological contributions that enable causal language model optimization on direct outcome datasets. First, we develop causal preference optimization (CPO), which solves an unbiased surrogate objective for the causal optimization problem. Next, we extend this to doubly robust CPO (DR-CPO), which improves on CPO by reducing the variance of the surrogate objective via outcome modeling while retaining provably strong guarantees on bias. [^1] Our code is publicly available at https://github.com/torylin/causal-preference-optimization.
@@ 3 A Causal View of Language Model Optimization
When a language model $f$ is trained or fine-tuned to generate texts that are consistent with human preferences, the implicit goal can be seen as optimizing texts $X\sim P^{f}$—texts generated from the model $f$—with respect to some outcome $Y$. In this paper, we consider a direct outcome data format $\mathcal{D}_{O}=\{(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\}$, where $X_{i}$ is a text that individual $i$ interacts with and $Y_{i}$ is any numerical response of the individual to the texts (e.g., ratings, either binary or scalar).
@@
These hypothetical outcomes can be formalized using the potential outcomes framework [Neyman, 1923 [1990], Rubin, 1974]: over text space $\mathcal{X}$, for each individual $i$ we posit the existence of a potential outcome function $Y_{i}:\mathcal{X}\to\mathbbm{R}$, where $Y_{i}(x)$ encodes their potential real-valued response if given text $x$ [^2] This notation implicitly rules out the possibility that an individual’s responses can be affected by the texts given to others—a common assumption in causal inference [Rubin, 1974]. We emphasize that most individuals’ potential outcomes are not observed, and so $Y_{i}(x)$ denotes the response individual $i$ would have had had they seen text $x$, possibly contrary to reality. This is also commonly known as the counterfactual.
@@
[…]
@@
Reframing our optimization problem as a causal inference problem allows us to draw on solutions from statistical causal inference—in particular, the use of randomized experiments to identify causal effects and approximate observing all potential outcomes. [^3] As discussed above, crowdsourced datasets are randomized experiments, since annotators are randomly assigned to read texts. We formalize such a randomized experiment and/or crowdsourced annotated dataset as $\mathcal{D}_{R}=\{(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\}$ where texts $X_{i}$ are drawn i.i.d. from a known randomization distribution $P^{R}$, and individuals with potential outcome functions $Y_{i}(\cdot)$ are drawn i.i.d. from the population $\mathcal{G}$ [^4] The probability of any text $X_{i}$ is known according to the randomization probability. For instance, if each reader reads one text uniformly assigned from a corpus of size $n$, then $P^{R}(X_{i})=\frac{1}{n}$ for all $X_{i}$ in the corpus. This induces a distribution on the observed responses $Y_{i}=Y_{i}(X_{i})$ that we denote as $P^{R}_{y}$.
@@ 4.1 Causal Preference Optimization
Note that both $P^{f}$ and $P^{R}$ are known quantities and do not need to be estimated—$P^{f}$ because it is obtained directly from the model $f$ we are optimizing, and $P^{R}$ because we know the randomization mechanism of the texts in $\mathcal{D}_{R}$. [^5] In practice, it can still be empirically helpful to use a model-derived estimate of the randomization probabilities $\widehat{P}^{R}(X)$, similar to how the Hájek estimator can have lower variance than the Horvitz-Thompson estimator [Hájek, 1971, Särndal et al., 2003]. Importantly, this means that $\widehat{V}_{IPW}(f)$ is an unbiased estimator for $V(f)$.
@@ 4.2 Doubly Robust Causal Preference Optimization
An importance weighting estimator like the CPO value function is a natural solution for estimating causal quantities. However, CPO optimizes over only randomized experimental data such as crowdsourced data, and so it can be further improved by the addition of an outcome modeling term that can leverage (often larger) non-randomized data to learn to predict outcomes on unlabeled texts. Doubly robust causal preference optimization (DR-CPO) combines IPW (over randomized data) and outcome modeling (over non-randomized data) to yield a doubly robust estimator that reduces the variance of CPO and improves its generality while still remaining unbiased for the true causal optimization problem.
@@
Importantly, because $P^{R}$ is known in randomized experiments, including the crowdsourced data setting, it does not need to be estimated, which means that condition (1) is always fulfilled for DR-CPO. As a result, $\widehat{V}_{DR}(f)$ is guaranteed to be unbiased for $V_{DR}$ even if the outcome model is incorrect. In other words, DR-CPO is robust to misspecification of $\widehat{g}$, as the IPW term in its value function corrects for any bias from the predicted outcomes. This means that rather than requiring a randomized dataset for training, $\widehat{g}$ can reasonably be learned over any direct outcome dataset, including those where the causal relationship between the text and the outcome may be confounded.
Therefore, in addition to learning from the experimental data $X\sim P^{R},Y\sim P^{R}_{y}$, a model optimized with DR-CPO will be able to further leverage the generative language model $f^{0}$ to learn from unlimited unlabeled text $\widetilde{X}\sim P^{f^{0}}$ with predicted outcomes $\widehat{g}(\widetilde{X})$. This can reduce the variance of the value function estimator.
@@
$\quad-\frac{n}{m}\text{Var}\left(\frac{P^{f}(\widetilde{X})}{P^{f^{0}}(\widetilde{X})}\widehat{g}(\widetilde{X})\right)$
@@
Condition (i) limits the component of the variance difference that arises due to Monte Carlo error from taking $m$ samples from the reference language model $f^{0}$. Then the main comparison is the difference between the variance of the expected outcome $g(x)$, and the variance of the prediction error for the model. Consequently, under condition (ii), we can expect the variance of $\widehat{V}_{DR}$ to be lower than the variance of $\widehat{V}_{IPW}$ (e.g., if $(\widehat{g}(x)-g(x))^{2}<g(x)^{2}$).
We note that in a different data setting where the true $P^{R}$ is unknown (e.g., no crowdsourced or randomized data is available), DR-CPO can still solve the true causal problem if all confounders are known in the dataset used to train $\widehat{g}$. Controlling for all confounders allows the outcome model to learn the correct causal relationship between the text and the outcome. Since such an outcome model is correct, it fulfills the condition $\widehat{g}(X)=g(X)$, and it follows from Theorem 4.4 that the DR-CPO objective $\widehat{V}_{DR}$ is an unbiased estimator for the true value function $V(f)$ even if $P^{R}$ is not randomized.
In practice, the full set of confounders is rarely known (particularly for text data). However, a good $\widehat{g}$ can still likely be learned if a large amount of clean data exists to train the outcome model. In this case, a well-estimated outcome model $\widehat{g}$ can still help bias-correct mis-estimation of $\widehat{P}^{R}(X)$.
@@ 4.3 Relationship to Existing Approaches
First, notice that the outcome modeling term $V_{out}$ is itself a way of identifying $V(f)$ from the observed data. Outcome modeling relies entirely on the predictive model $\widehat{g}$ and unlabeled texts $\widetilde{X}\sim P^{f^{0}}$ and therefore does not require an experimental dataset $\mathcal{D}_{R}$. However, if $\widehat{g}(X)$ is not a good outcome model, then $\widehat{V}_{out}(f)$ will be biased with respect to $V(f)$. Reward models trained on confounded data may be misspecified, since they can only capture the relationship between the text and the response—and not the confounders that have additionally influenced the response. While these issues are remedied if the confounding is also fully modeled, confounders are extremely difficult to measure fully in text data.
Optimization of $V_{out}(f)$ is closely related to RLHF in the direct outcome setting via proximal policy optimization (PPO) and REINFORCE [Stiennon et al., 2020, Williams, 1992, Ahmadian et al., 2024]. In particular, the PPO loss term can be seen as a version of outcome modeling in which the reward model is trained on paired completion data. Under these conditions, the RLHF reward model is analogous to the outcome model $g$, and the PPO loss is mathematically equivalent to $V_{out}(f)$ (details in Appendix B). Because in practice PPO often optimizes both the policy network and the value function network, $V_{out}(f)$—which optimizes only the policy network—can also be seen as similar to a REINFORCE objective in which learning occurs over Monte Carlo samples drawn from a fixed language model $f^{0}$.
Likewise, DPO shares similarities with CPO. Rather than relying on reward or outcome modeling, both DPO and CPO fine-tune a language model for human feedback by directly using a preference dataset, which greatly reduces their computational overhead. The DPO objective is similar to $V_{IPW}$ in that it directly increases the likelihood of texts corresponding to desired outcomes through importance weighting. However, because it uses paired data, DPO increases the density ratio between preferred and non-preferred examples, while CPO directly increases the probability of texts with desired outcomes and decreases the probability of texts with non-desired outcomes. Given a paired data setting, the DPO objective could possibly be recovered from $V_{IPW}$; we leave this derivation for future work.
@@ 5.1 Datasets
Hate Speech (binary outcome). The Hate Speech dataset [Qian et al., 2019] consists of comments from the social media sites Reddit and Gab. Comments span a wide variety of subjects, including daily life, personal relationships, politics, and finance. Outcomes are collected via crowdsourcing and indicate whether the annotator perceives the comment to be hate speech. The Reddit comments are chosen from subreddits where hate speech is more common, and Gab is a platform where users sometimes migrate after being blocked from other social media sites. The optimization goal for this dataset is to generate texts that are less hateful on average.
We include this dataset to evaluate optimization under conditions where (i) texts are topically diverse and (ii) outcomes are relatively simple to judge.
@@
We include this dataset to evaluate optimization under more challenging conditions where (i) texts all concern a single topic and (ii) outcomes are complex, noisy, and difficult for a model to predict.
Confounded (scalar outcome). The Confounded dataset is a version of the Hong Kong dataset where we have induced confounding. Concretely, it is identical to the Hong Kong dataset with the exception of its outcome, which is the negation of the original outcome (e.g., an outcome of 60 in the Hong Kong dataset becomes an outcome of -60 in the Confounded dataset). This inverts the relationship between the text and the outcome: texts originally associated with high outcomes in the Hong Kong dataset are now associated with low outcomes, and vice versa. Intuitively, negating the outcome is akin to introducing a strong confounding or corrupting factor where readers who tend to read texts that should have high outcomes also tend to have very negative responses to those texts (e.g., “haters” brigading a popular celebrity’s posts).
We include this dataset with the realistic expectation that text data is often confounded. While CPO by construction uses only randomized data, confounded data can pose a threat to approaches that use non-randomized data to train their outcome models. Therefore, it is necessary to evaluate how approaches like DR-CPO and OO-RLHF fare when their outcome models are trained on confounded data.
@@ 5.2 Implementation
Evaluation. To evaluate how well optimization for human preferences has occurred, we consider two separate metrics. As our primary evaluation, we use a text preference framework in which a reader is asked to choose the better (with respect to the outcome) of a pair of texts generated by two different methods. Using GPT-4 as a proxy for human annotators, we compare pairs of (method, baseline) completions for the same prompt; across all pairs, we compute method win rates and compute 95% confidence intervals. Since the datasets used for these experiments contain one text per sample rather than a prompt and a completion, we create prompts on the evaluation set by truncating each text to a random length.
@@
To supplement our text preference evaluation, we additionally assess the effectiveness of each method in maximizing expected reward (i.e., the value function). In particular, we focus on the reward under the DR-CPO objective, which is the most efficient estimate of the reward.
Methods. We evaluate language models optimized using CPO and DR-CPO. As our baselines, we consider language models that have been fine-tuned on texts from each of the task datasets (FT), as well as models optimized using the outcome modeling value function $V_{out}(f)$. Since—as we discuss in Section 4.3—the $V_{out}(f)$ objective is mathematically equivalent to the RLHF objective, we refer to this baseline as OO-RLHF (offline outcome RLHF).
We use Llama 2 7B [Touvron et al., 2023] as our base language model and fine-tune with low-rank adaptation (LoRA) [Hu et al., 2022]. All optimizations are applied after fine-tuning on text from the task dataset. Because only a tiny fraction of parameters are updated under LoRA, we note that we expect relatively small changes in performance under optimization.
Our outcome models for the Hate Speech dataset are also based on Llama 2 7B. For the Hong Kong outcome model, LLMs perform poorly on the dataset’s complex outcomes, particularly given the dataset’s smaller size. Instead, we use a standard linear regression model over features extracted using the Empath lexicon [Fast et al., 2016].
Implementation details can be found in Appendix C.1.
Choice of $f^{0}$. When optimizing with DR-CPO or OO-RLHF, any generative language model may be used as $f^{0}$, the fixed language model from which texts are drawn as input to the outcome model. One key consideration is whether $f^{0}$ should be a pre-trained model or whether it should be a model that has been fine-tuned on text relevant to the task—for instance, the randomized experiment dataset $\mathcal{D}_{R}$.
In practice, we choose a pre-trained model as $f^{0}$ to leverage the diversity of texts such models tend to generate. If $\widehat{g}(X)$ is a good outcome model, then predicted outcomes on these texts will still be close to the true outcomes, and DR-CPO and OO-RLHF will benefit from outcome modeling.
Choice of $\widehat{P}^{R}$. In Section 4.1, we mention that although the distribution of texts $P^{R}(X)$ under the randomized experiment is known, it can be helpful empirically to instead compute an estimated $\widehat{P}^{R}(X)$. This is generally due to the fact that the sample probability of each text $X$ may not actually be equal to its theoretical probability merely by chance [Hájek, 1971, Särndal et al., 2003].